# Optimal quantum communication networks: capacitance versus security

### v1

@@ Abstract
The rate and security of quantum communications between users placed at arbitrary points of a quantum communication network depend on the structure of the network, on its extension and on the nature of the communication channels. In this work we propose a strategy of network optimization that intertwines classical network approaches and quantum information theory. Specifically, by suitably defining a quantum efficiency functional, we identify the optimal quantum communication connections through the network by balancing security and the quantum communication rate. The optimized network is then constructed as the network of the maximal quantum efficiency connections and its performance is evaluated by studying the scaling of average properties as functions of the number of nodes and of the network spatial extension.
@@ I Introduction
Quantum communication networks [1] enable the realization of tasks beyond the reach of classical communication systems. Examples are unconditionally secure quantum key distribution [2, 3] (QKD), quantum teleportation [4], clock-synchronization [5], distributed quantum computing [6], to mention just a few. Characterizing and optimizing quantum communication networks have both a crucial relevance for the development of quantum cryptography applications [7] and for the potential to advance our understanding of fundamental quantum phenomena [8], such as entanglement percolation [9] or the emergence of non-local quantum correlations[10, 11, 12].
@@
In the ideal case the ultimate properties of network elements, such as the quantum communication links, are dictated by the laws of quantum mechanics, enforcing their security but also imposing intrinsic bounds [26] on the rate of quantum information transmission. A particularly relevant example is the Pirandola, Laurenza, Ottaviani and Banchi [27] (PLOB) bound that prevents to achieve simultaneously high rates and long distances in transferring quantum states and distributing entanglement or secret quantum keys through a quantum link.
@@
In this work we employ the tools of classical network science to devise a strategy of optimization of quantum communication networks. PLOB bound can be indeed circumvented by means of intermediate repeaters [13], either of quantum [30] or classical nature (trusted nodes [31, 32]), that help the communication between distant parties. With few notable exceptions [33], most field tests of metropolitan-scale quantum networks to date are based on point-to-point architecture and they involve trusted nodes, see Refs. [34, 35, 36, 37, 38]. Trusted nodes in general lower the security of the network [39, 31, 40]. Consequently, in designing a QKD network the following question arises naturally: given a set of QKD users, what is the optimal way to connect them to fulfill a given rate/security target, assuming that all trusted nodes have a certain probability $p$ of being leaky? Here we address this optimization problem by introducing a quantum efficiency functional, that balances the quantum security and the quantum communication rate for each pair of users in the network. We develop an algorithm that maximizes the quantum efficiency and constructs the optimal network, that we refer to as maximal quantum efficiency network. We then investigate, for a random distribution of users in the plane, the average properties of these optimal networks. Their performance is evaluated by studying the scaling of average properties as a function of the number of nodes and of the network extension. While the quantum communication rate is linked to geometrical properties, such as the average distance between users, the security depends on the topology. The optimization algorithm therefore goes beyond standard dynamic programming methods such the Dijkstra algorithm [41], that was previously employed in the context of quantum repeater network optimization [21].
@@ II Results
II.1 Efficiency of quantum networks
@@
Within the model defined by Eqs. (2) and (3), capacitance and security in general compete, i.e., longer paths may have larger capacitance but at the price of lower security. To describe this trade-off we define the efficiency $\epsilon_{\alpha}(\{a\to b\})$ of a path
@@
where $\alpha\in[0,1]$ is a user-tunable control parameter which gives more importance either to capacitance ($\alpha=0$) or to security ($\alpha=1$). For simplicity the efficiency is defined using the logarithm of the security that is proportional to the path length.
Given the positions of users in space and assuming that a direct link can be placed between any two of them, so that the network of all possible links ${\cal G}_{0}$ is a fully connected (FC) graph, there is an exponentially large number of possible paths between two nodes, $a$ and $b$. For a given $\alpha$, an optimal path $\{a\to b\}^{*}$ is defined as a path having maximal efficiency among all possible paths
@@
Clearly, in the limit $\alpha\to 0$, the optimal path maximizes the capacitance, by going through many, physically close, intermediate nodes. In the limit $\alpha\to 1$ the optimal path maximizes security and is thus the direct link between the two nodes. The global efficiency of a network ${\cal G}$ is defined as the average, over all node pairs, of the efficiency of the optimal paths defined over the network
@@
It is important to remark that in general there can be more than a single optimal path between two nodes. As an example in Fig. 1(a) we show two optimal paths between the nodes $a$ and $b$ passing through the node $k$ and $k^{\prime}$, respectively. The two paths have the same efficiency since they have the same topological length, $\ell=4$, and the same capacitance $Q=q(e_{cd})$, being $e_{cd}$ the longest edge in both paths.
@@ II.2 Network optimization
Given a set of $N$ users, ${\cal N}$, distributed in a square of size $L$, and a tradeoff parameter $\alpha$, our goal is to find a network connecting the $N$ users having maximal efficiency and minimal number of links. To reach this goal we start by determining an optimal path for each pair of users. This is achieved by means of a novel algorithm described in detail in the Methods section. Once an optimal path is identified for each pair, we define the optimal network, ${\cal G}^{*}_{\alpha}\equiv({\cal N},{\cal E}_{\alpha}^{*})$, as the graph union of the optimal paths, that is the network with node set ${\cal N}$ and edge set ${\cal E}_{\alpha}^{*}$ given by
@@
This network has the maximum quantum efficiency, i.e. no other network ${\cal G^{\prime}}\equiv({\cal N},{\cal E}^{\prime})$ can have larger efficiency, i.e.,
@@
The method developed in this work builds on Pollack’s algorithm to construct the optimal network $\cal G_{\alpha}^{*}$ as prescribed by Eq.(7). Figure 1(c-f) shows instances of $\cal G_{\alpha}^{*}$ for different values of $\alpha$. As one can see, it interpolates between the fully connected network, realized for $\alpha=1$, and a much sparser network for $\alpha\to 0$. We note that for $\alpha\to 0$, $\cal G_{\alpha}^{*}$ does not necessarily reduce to the MST. This is related to the fact that, as we explained above, the optimal paths are non-unique. In principle, for generic $\alpha$ one could complement our algorithm with further optimization techniques to reduce the number of links of $\cal{G}_{\alpha}^{*}$ while preserving maximum efficiency. This further development is deferred to future work.
II.3 Maximal quantum efficiency networks
Figure 2: Average capacitance and path length, $Q^{*}$ and $\mathcal{L}^{*}$, on the MQE networks as a function of $\alpha$ for $p=1-1/e\simeq 0.63$. $N=256$ (circles), $N=512$ (triangles), $N=1024$ (diamonds), $N=2048$ (squares). a) $L/\lambda_{0}=0.1$, b) $L/\lambda_{0}=1$, c) $L/\lambda_{0}=3$, d) $L/\lambda_{0}=10$. Averages over $100$ realizations for $N=256$ and $N=512$, and over $10$ realizations for $N=1024$ and $N=2048$. Dashed lines correspond to the values of $\alpha_{c}^{0\to 1}$ (red), $\alpha_{c}^{1\to 2}$ (green) and $\alpha_{c}^{2\to 3}$ (purple) evaluated using Eq. (12).
For concreteness, in this subsection we assume a completely random distribution of users in the square of size $L$. As highlighted above, optimal networks maximize the quantum efficiency functional defined by Eq. (6). Their structure and performance depend sensitively on the user-defined parameter $\alpha$ and on the distance-loss ratio, $L/\lambda_{0}$, that set the optimization regime. To show how these maximal quantum efficiency (MQE) networks change across the different regimes, in Fig. 2 we plot their average capacitance,
@@ II.3 Maximal quantum efficiency networks
(i) For values of $\alpha$ larger than a threshold value $\alpha_{c}$ we observe *${\mathcal{L}^{*}}=1$, corresponding to all optimal paths having $\ell_{\{a\to b\}^{*}}=1$. In this regime ${\cal G}_{\alpha}^{*}$ coincides with the fully connected network. Moreover, since the structure of ${\cal G}_{\alpha}^{*}$ does not change with $\alpha$, the average capacitance is constant, $Q^{*}=Q_{\text{FC}}$ (see Methods for its evaluation). The critical value $\alpha_{c}$ is the lowest value of $\alpha$ such that, for every pair of users $a,b$ in the system, it is more efficient to connect them through the direct link rather than using a trusted node $c$. This happens as long as the gain in efficiency due to increased capacitance, is smaller than the loss of efficiency associated to the introduction of an intermediate node, equal to $\alpha\log(1-p)$, yielding (see Methods for details)
@@
(ii) On the left of this threshold value we observe the existence of a “step”, i.e., an interval of $\alpha$ values over which $\mathcal{L}^{*}$ is practically constant and equal to $2$. In this interval all optimal paths have 1 intermediate node. As $\alpha$ decreases other steps appear, corresponding to MQE networks featuring all optimal paths having topological length *${\cal L}^{*}=3$ (second step) and *${\cal L}^{*}=4$ (third step). The steps become sharper in the large $N$ limit where a simple calculation (see Methods) shows that paths having $m$ intermediate nodes become more efficient than those with $m-1$ intermediate ones for
@@
(iii) As $\alpha$ tends to 0, MQE networks increasingly resemble the maximum spanning tree. In this limit the average topological length of optimal paths tends to increase as a power-law with $N$ (see Methods). Let us now consider the opposite case of strong losses, $L\gg\lambda_{0}$. In this limit the phenomenology is different and more difficult to interpret because, while most distances are much larger than $\lambda_{0}$, still some pairs of users are at a distance smaller than the decay length. These pairs are responsible for the observation that *$\mathcal{L}^{*}>1$ as soon as $\alpha$ gets smaller than the threshold value still given by Eq. (11) (see Fig. 2(c-d)). At variance with the previous case, for $\alpha<\alpha_{c}$ the average topological length of the optimal paths $\mathcal{L}^{*}$ grows rapidly as $\alpha$ is reduced, showing no steps at integer values. Moreover, there exist a value $\overline{\alpha}$ such that for $\alpha<\overline{\alpha}$, $\mathcal{L}^{*}$ and $Q^{*}$ vary much less and assume values close to those of the MST network (see Methods).
@@
Figure 3: (a) Transitions between the different regimes for $p=1-1/e$ and $N\to\infty$, representing the number of intermediate nodes in the optimal path for a pair of nodes at distance $d_{ab}$ as a function of $d_{ab}$ and $\alpha$. Note that the curve separating the “direct link region” from the other regions has slope discontinuities corresponding to the dashed vertical lines. (b) Behavior of the average capacitance $Q^{*}$, and of the minimum capacitance $Q_{\rm min}$ as a function of $L/\lambda_{0}$ with $\alpha=0.1$, $p=0.1$, $N=100$ for the MQE network and for the FC network (computed analytically, see Methods). (c) Plot of the average capacitance of the MQE network as a function of the average topological length for $p=1-1/e$, $N=512$ (dotted lines), $N=1024$ (solid lines) and various values of $L/\lambda_{0}$.
@@
II.3.1 Performance of MQE networks
By balancing security and quantum capacitance, MQE networks yield the optimal strategy to connect $N$ users for quantum communications and, at the same time, they represent a customizable benchmarking tool for quantum communication networks. To illustrate the performance of MQE networks we start by comparing their quantum capacitance to that of FC networks. As one can see in Fig. 3(b), the outcome depends sensitively on the ratio $L/\lambda_{0}$. For $L/\lambda_{0}\ll 1$ a limited increase of the average capacitance with respect to the FC network is observed. For values of $L$ of the order of $\lambda_{0}$ the improved performance of the MQE network becomes more significant. In this range of $L/\lambda_{0}$ values, the most important improvement concerns the minimum capacitance. While for FC networks some links have a strongly degraded capacitance, several orders of magnitude smaller than the average $Q^{*}$, in the optimized network $\min Q$ is only slightly smaller than the average, thus guaranteeing that communication is possible among any pairs of users. For very large values of $L/\lambda_{0}$ the MQE network tends to coincide with the FC one, as the weight of capacitance in the quantum efficiency functional becomes extremely small. In this regime the average capacitance $Q^{*}$ is essentially the same in the two networks and decays as $(L/\lambda_{0})^{-2}$. In Fig. 3(b) we also note that in MQE networks the minimum capacity reaches the threshold of one target bit per use of the channel, $Q_{\rm min}=1$, for values of $L/\lambda_{0}$ about one order of magnitude larger than in standard FC networks. For fixed values of $\lambda_{0}$, this implies that MQE networks can provide the minimum quantum communication rate standard across much wider regions of space. Specifically, for the parameters in Fig. 3(b), the maximum value of $L$ to have at least 1 target bit transferred between any user of the network goes from $\approx 0.5\lambda_{0}$ (in the FC network) to $\approx 5\lambda_{0}$ in the MQE network. A natural question arises concerning how the the security requirement affects the maximal achievable rate. To clarify this point in Fig. 3(c) we plot the average capacitance of MQE networks, $Q^{*}$ as a function of the average topological length ${\cal L}^{*}$, for different values of the ratio $L/\lambda_{0}$. It turns out that it is sufficient to allow $\cal L^{*}$ to grow from 1 (FC network) to 3 or 4 to ensure a considerable increase in the average capacitance, even one order of magnitude for large $L/\lambda_{0}$.
II.3.2 Structure of MQE networks

### v3

# Optimal quantum key distribution networks: capacitance versus security

@@ Abstract
The rate and security of quantum communications between users placed at arbitrary points of a quantum communication network depend on the structure of the network, on its extension and on the nature of the communication channels. In this work we propose a strategy for the optimization of trusted-relays based networks that intertwines classical network approaches and quantum information theory. Specifically, by suitably defining a quantum communication efficiency functional, we identify the optimal quantum communication connections through the network by balancing security and the quantum communication rate. The optimized network is then constructed as the network of the maximal quantum communication efficiency connections and its performance is evaluated by studying the scaling of average properties as functions of the number of nodes and of the network spatial extension.
@@ I Introduction
Quantum communication networks [1] enable the realization of tasks beyond the reach of classical communication systems. Examples are unconditionally secure quantum key distribution [2, 3] (QKD), quantum teleportation [4], clock-synchronization [5], distributed quantum computing [6], to mention just a few. Characterizing and optimizing quantum communication networks have a crucial relevance for the development of quantum cryptography applications [7] and hold the potential to advance our understanding of fundamental quantum phenomena [8], such as entanglement percolation [9] or the emergence of non-local quantum correlations[10, 11, 12].
@@
In the ideal case the ultimate properties of network elements, such as the quantum communication links, are dictated by the laws of quantum mechanics, enforcing their security but also imposing intrinsic bounds [26, 27] on the rate of quantum information transmission. Specifically, the fundamental limit of repeaterless quantum communication found by Pirandola, Laurenza, Ottaviani and Banchi [28], known as PLOB bound, prevents to achieve simultaneously high rates and long distances in transferring quantum states and distributing entanglement or secret quantum keys through a quantum link.
@@
In this work we employ the tools of classical network science to devise a strategy of optimization of quantum communication networks. PLOB bound can be indeed circumvented by means of intermediate repeaters [13], either of quantum [31] or classical nature (trusted nodes [32, 33]), that help the communication between distant parties. With few notable exceptions [34], most field tests of metropolitan-scale quantum networks to date are based on point-to-point architecture and they involve trusted nodes, see Refs. [35, 36, 37, 38, 39]. Trusted nodes in general lower the security of the network [40, 32, 41]. Consequently, in designing a QKD network the following question arises naturally: given a set of QKD users, what is the optimal way to connect them to fulfill a given rate/security target, assuming that all trusted nodes have a certain probability $p$ of being leaky? Here we address this optimization problem by introducing a quantum communication efficiency functional, that balances the quantum security and the quantum communication rate for each pair of users in the network. Note that in a classical network leakage can occur not only at nodes where the signal is amplified, but also along the connections between nodes: adding amplifiers therefore improves the capacitance without necessarily reducing the security level. For this reason in classical networks there is no tradeoff between capacitance and security, which is instead inherent in QKD networks. We develop an algorithm that maximizes the quantum communication efficiency and constructs the optimal network, that we refer to as maximal quantum communication efficiency network. We then investigate, for a random distribution of users in the plane, the average properties of these optimal networks. Their performance is evaluated by studying the scaling of average properties as a function of the number of nodes and of the network extension. While the quantum communication rate is linked to geometrical properties, such as the average distance between users, the security depends on the topology. The optimization algorithm therefore goes beyond standard dynamic programming methods, such as the Dijkstra algorithm [42], that were previously employed in the context of quantum repeater network optimization [21].
@@ II Results
II.1 Communication efficiency of quantum networks
@@
Within the model defined by Eqs. (2) and (3), capacitance and security in general compete, i.e., longer paths may have larger capacitance but at the price of lower security. To describe this trade-off we define the communication efficiency $\epsilon_{\alpha}(\{a\to b\})$ of a path
@@
where $\alpha\in[0,1]$ is a user-tunable control parameter which gives more importance either to capacitance ($\alpha=0$) or to security ($\alpha=1$). For simplicity the communication efficiency is defined using the logarithm of the security that is proportional to the path length.
Given the positions of users in space and assuming that a direct link can be placed between any two of them, so that the network of all possible links ${\cal G}_{0}$ is a fully connected (FC) graph, there is an exponentially large number of possible paths between two nodes, $a$ and $b$. For a given $\alpha$, an optimal path $\{a\to b\}^{*}$ is defined as a path having maximal communication efficiency among all possible paths
@@
Clearly, in the limit $\alpha\to 0$, the optimal path maximizes the capacitance, by going through many, physically close, intermediate nodes. In the limit $\alpha\to 1$ the optimal path maximizes security and is thus the direct link between the two nodes. The global communication efficiency of a network ${\cal G}$ is defined as the average, over all node pairs, of the communication efficiency of the optimal paths defined over the network
@@
It is important to remark that in general there can be more than a single optimal path between two nodes. As an example in Fig. 1(a) we show two optimal paths between the nodes $a$ and $b$ passing through the node $k$ and $k^{\prime}$, respectively. The two paths have the same communication efficiency since they have the same topological length, $\ell=4$, and the same capacitance $Q=q(e_{cd})$, being $e_{cd}$ the longest edge in both paths.
@@ II.2 Network optimization
Given a set of $N$ users, ${\cal N}$, distributed in a square of size $L$, and a tradeoff parameter $\alpha$, our goal is to find a network connecting the $N$ users having maximal communication efficiency and minimal number of links. To reach this goal we start by determining an optimal path for each pair of users. This is achieved by means of a novel algorithm described in detail in the Methods section. Once an optimal path is identified for each pair, we define the optimal network, ${\cal G}^{*}_{\alpha}\equiv({\cal N},{\cal E}_{\alpha}^{*})$, as the graph union of the optimal paths, that is the network with node set ${\cal N}$ and edge set ${\cal E}_{\alpha}^{*}$ given by
@@
This network has the maximum quantum communication efficiency, i.e. no other network ${\cal G^{\prime}}\equiv({\cal N},{\cal E}^{\prime})$ can have larger communication efficiency, i.e.,
@@
The method developed in this work builds on Pollack’s algorithm to construct the optimal network $\cal G_{\alpha}^{*}$ as prescribed by Eq.(7). Figure 1(c-f) shows instances of $\cal G_{\alpha}^{*}$ for different values of $\alpha$. As one can see, it interpolates between the fully connected network, realized for $\alpha=1$, and a much sparser network for $\alpha\to 0$. We note that for $\alpha\to 0$, $\cal G_{\alpha}^{*}$ does not necessarily reduce to the MST. This is related to the fact that, as we explained above, the optimal paths are non-unique. In principle, for generic $\alpha$ one could complement our algorithm with further optimization techniques to reduce the number of links of $\cal{G}_{\alpha}^{*}$ while preserving maximum communication efficiency. This further development is deferred to future work.
II.3 Maximal quantum communication efficiency networks
Figure 2: Average properties of optimal networks. Average capacitance and path length, $Q^{*}$ and $\mathcal{L}^{*}$, on the MQCE networks as a function of $\alpha$ for $p=1-1/e\simeq 0.63$. $N=256$ (circles), $N=512$ (triangles), $N=1024$ (diamonds), $N=2048$ (squares). a) $L/\lambda_{0}=0.1$, b) $L/\lambda_{0}=1$, c) $L/\lambda_{0}=3$, d) $L/\lambda_{0}=10$. Averages over $100$ realizations for $N=256$ and $N=512$, and over $10$ realizations for $N=1024$ and $N=2048$. Dashed lines correspond to the values of $\alpha_{c}^{0\to 1}$ (red), $\alpha_{c}^{1\to 2}$ (green) and $\alpha_{c}^{2\to 3}$ (purple) evaluated using Eq. (12).
For concreteness, in this subsection we assume a completely random distribution of users in the square of size $L$. As highlighted above, optimal networks maximize the quantum communication efficiency functional defined by Eq. (6). Their structure and performance depend sensitively on the user-defined parameter $\alpha$ and on the distance-loss ratio, $L/\lambda_{0}$, that set the optimization regime. To show how these maximal quantum communication efficiency ( MQCE) networks change across the different regimes, in Fig. 2 we plot their average capacitance,
@@ II.3 Maximal quantum communication efficiency networks
(i) For values of $\alpha$ larger than a threshold value $\alpha_{c}$ we observe ${\mathcal{L}^{*}}=1$, corresponding to all optimal paths having $\ell_{\{a\to b\}^{*}}=1$. In this regime ${\cal G}_{\alpha}^{*}$ coincides with the fully connected network. Moreover, since the structure of ${\cal G}_{\alpha}^{*}$ does not change with $\alpha$, the average capacitance is constant, $Q^{*}=Q_{\text{FC}}$ (see Methods for its evaluation). The critical value $\alpha_{c}$ is the lowest value of $\alpha$ such that, for every pair of users $a,b$ in the system, it is more efficient to connect them through the direct link rather than using a trusted node $c$. This happens as long as the gain in communication efficiency due to increased capacitance, is smaller than the loss of communication efficiency associated to the introduction of an intermediate node, equal to $\alpha\log(1-p)$, yielding (see Methods for details)
@@
(ii) On the left of this threshold value we observe the existence of a “step”, i.e., an interval of $\alpha$ values over which $\mathcal{L}^{*}$ is practically constant and equal to $2$. In this interval all optimal paths have 1 intermediate node. As $\alpha$ decreases other steps appear, corresponding to MQCE networks featuring all optimal paths having topological length ${\cal L}^{*}=3$ (second step) and ${\cal L}^{*}=4$ (third step). The steps become sharper in the large $N$ limit where a simple calculation (see Methods) shows that paths having $m$ intermediate nodes become more efficient than those with $m-1$ intermediate ones for
@@
(iii) As $\alpha$ tends to 0, MQCE networks increasingly resemble the maximum spanning tree. In this limit the average topological length of optimal paths tends to increase as a power-law with $N$ (see Methods). Let us now consider the opposite case of strong losses, $L\gg\lambda_{0}$. In this limit the phenomenology is different and more difficult to interpret because, while most distances are much larger than $\lambda_{0}$, still some pairs of users are at a distance smaller than the decay length. These pairs are responsible for the observation that $\mathcal{L}^{*}>1$ as soon as $\alpha$ gets smaller than the threshold value still given by Eq. (11) (see Fig. 2(c-d)). At variance with the previous case, for $\alpha<\alpha_{c}$ the average topological length of the optimal paths $\mathcal{L}^{*}$ grows rapidly as $\alpha$ is reduced, showing no steps at integer values. Moreover, there exist a value $\overline{\alpha}$ such that for $\alpha<\overline{\alpha}$, $\mathcal{L}^{*}$ and $Q^{*}$ vary much less and assume values close to those of the MST network (see Methods).
@@
Figure 3: The phase diagram and the behavior of the average capacitance. (a) Transitions between the different regimes for $p=1-1/e$ and $N\to\infty$, representing the number of intermediate nodes in the optimal path for a pair of nodes at distance $d_{ab}$ as a function of $d_{ab}$ and $\alpha$. Note that the curve separating the “direct link region” from the other regions has slope discontinuities corresponding to the dashed vertical lines. (b) Behavior of the average capacitance $Q^{*}$, and of the minimum capacitance $Q_{\rm min}$ as a function of $L/\lambda_{0}$ with $\alpha=0.1$, $p=0.1$, $N=100$ for the MQCE network , the FC network (computed analytically, see Methods). (c) Plot of the average capacitance of the MQCE network as a function of the average topological length for $p=1-1/e$, $N=512$ (dotted lines), $N=1024$ (solid lines) and various values of $L/\lambda_{0}$.
@@
II.3.1 Performance of MQCE networks
By balancing security and quantum capacitance, MQCE networks yield the optimal strategy to connect $N$ users for quantum communications and, at the same time, they represent a customizable benchmarking tool for quantum communication networks. To illustrate the performance of MQCE networks we start by comparing their quantum capacitance to that of FC networks. As one can see in Fig. 3(b), the outcome depends sensitively on the ratio $L/\lambda_{0}$. For $L/\lambda_{0}\ll 1$ a limited increase of the average capacitance with respect to the FC network is observed. For values of $L$ of the order of $\lambda_{0}$ the improved performance of the MQCE network becomes more significant. In this range of $L/\lambda_{0}$ values, the most important improvement concerns the minimum capacitance. While for FC networks some links have a strongly degraded capacitance, several orders of magnitude smaller than the average $Q^{*}$, in the optimized network $\min Q$ is only slightly smaller than the average, thus guaranteeing that communication is possible among any pairs of users. For very large values of $L/\lambda_{0}$ the MQCE network tends to coincide with the FC one, as the weight of capacitance in the quantum communication efficiency functional becomes extremely small. In this regime the average capacitance $Q^{*}$ is essentially the same in the two networks and decays as $(L/\lambda_{0})^{-2}$. In Fig. 3(b) we also note that in MQCE networks the minimum capacity reaches the threshold of one target bit per use of the channel, $Q_{\rm min}=1$, for values of $L/\lambda_{0}$ about one order of magnitude larger than in standard FC networks. For fixed values of $\lambda_{0}$, this implies that MQCE networks can provide the minimum quantum communication rate standard across much wider regions of space. Specifically, for the parameters in Fig. 3(b), the maximum value of $L$ to have at least 1 target bit transferred between any user of the network goes from $\approx 0.5\lambda_{0}$ (in the FC network) to $\approx 5\lambda_{0}$ in the MQCE network. A natural question arises concerning how the the security requirement affects the maximal achievable rate. To clarify this point in Fig. 3(c) we plot the average capacitance of MQCE networks, $Q^{*}$ as a function of the average topological length ${\cal L}^{*}$, for different values of the ratio $L/\lambda_{0}$. It turns out that it is sufficient to allow $\cal L^{*}$ to grow from 1 (FC network) to 3 or 4 to ensure a considerable increase in the average capacitance, even one order of magnitude for large $L/\lambda_{0}$.
II.3.2 Structure of MQCE networks